# Neural Correlates of Language Models Are Specific to Human Language

### v1

@@ 3 Methodology
We used the dataset provided by pereira2018toward to compare brain and language model representations. This is consistent with previous studies (schrimpfNeuralArchitectureLanguage2021a, ; gauthierLinkingArtificialHuman2019, ; schwartzInducingBrainrelevantBias2019, ; ootaNeuralLanguageTaskonomy2022, ; awInstructiontuningAlignsLLMs2024, ) [^1] Dataset available at https://osf.io/crwz7/.. For the brain model comparisons, we used the fMRI data from 6 adult subjects. The stimuli sentences consisted of simple and natural statements ($N_{\text{stimuli}}=243$), with an average length of 13 words ($\sigma=2.9$). The subjects were indicated to read the sentences naturally and thinking about their meaning. The brain activity was recorded while subjects were presented the sentences one at a time. For each subject, the dataset provides 243 BOLD signals represented as $\approx$200,000-dimensional vectors.
@@ 3.4 Topological Analysis of Brain and LM Layers
In addition to GW, we also computed the centered kernel alignment (CKA) of each participant to each layer, as well as the overall mean for the [+pos] and [-pos] conditions. CKA allowed to measure how similarly two sets of representations encoded relationships across points with minimal intermediate steps. CKA has been previously used in the context of cognitive science cohen1994localization , deep learning kornblith2019similarity and alignment of artificial and biological neural networks murphy2024correcting . We followed the unbiased kernel alignment formulation kornblith2019similarity . The selection of this variant was motivated by previous findings on the limitations of the original CKA implementation (e.g., murphy2024correcting ). The selected variant has proven robust to high-sample low-dimensionality and low-sample high-dimensionality data setups (murphy2024correcting, ).
@@ 4.1 Positional Information is Key for Brain Alignment
Figure 1: Average brain scores (brain-model representation correlations) per condition (PCA and original dimensions; with and without positional encodings). These results were obtained by averaging all the results for all the layers of bidirectional (left) and causal (right) models. This shows a high level picture of the data.
Figure 2 shows each model’s layer-wise brain alignment as a function of relative layer position.
Layer trajectories in bidirectional models. Patterns diverged across architectures: some displayed a smooth monotonic rising (e.g., distilbert-uncased), whereas others dipped in the middle layers before recovering at the deepest (albert-large-v2). Several models maintained above-average [–pos] scores, pointing to implicit positional inference. All bidirectional encoders peaked in the final layers.
@@
Figure 2: All models’ layer-wise brain scores (left bidirectional models, right causal models). The results of the [+pos] condition are shown in blue; orange shows the results for [-pos] condition. Dashed lines show results using PCA data. Purple line indicates overall mean. 0% indicates input layer; 100% corresponds to output layer.
@@ 4.2 Scaled Dot-product Attention Accounts for Some Brain Alignment
To understand how models’ internal mechanisms behaved in the observed brain alignment and experimental conditions, we analyzed the scaled dot-product attention heads’ brain scores. Analyzing individualized contributions helped clarify whether the observed brain alignment showed distributed properties or specialization clusters within the attention mechanism. Head-wise brain scores were computed as in the layer-wise analysis (§3.3 and §4.1). Figure 3 reports scores for the [-pos] and [+pos] conditions.
@@
Figure 3: Head-wise (scaled dot-product attention) brain scores for bidirectional (top) and causal (bottom) models. Brighter colors indicate better correlations. Lengths vary depending of the model size: bigger models have more heads.
Figure 4: Head-wise (attention weights only) brain scores for bidirectional (top) and causal (bottom) models. Brighter colors indicate better correlations. Lengths vary depending of the model size: bigger models are composed by more attention heads.
@@ 4.3 Attention Weights Are Marginally Relevant for Brain Alignment
Bidirectional encoders. Across all bidirectional models, attention weights’ brain scores remained uniformly low in the [–pos] and recovered some relevance at the [+pos] condition (Figures 4 and 3). However, adding positional encodings produced only marginal gains, far smaller than those observed at layer‑level or scaled dot-product attention analyses. Moreover, the changes did not exhibit a coherent depth‑wise trend: correlations fluctuated idiosyncratically from head to head, suggesting that no specific subset of bidirectional heads was consistently responsible for brain alignment.
@@ 4.4 Brain and Layer Data Spaces Share Characteristics
Layer-wise CKA Analysis. All participants’ brain representations were severely misaligned when the model was deprived of positional information. Interestingly, most subjects showed a similar pattern: CKA steadily decreased to the 0.25-0.37 range (layer 20 to 25) showing a momentous recovery at the 28th layer, and a final decrease from there to the last layer. Adding positional information drastically changed the trends, showing a largely monotonic improvement until reaching the peak similarities ($\approx$0.61-0.87) at the final layers. The results shown in Figure 5 complement those from the previous experiments.
Figure 5: Participant-model CKA results for [-pos] (left) and [+pos] (right) conditions. Each line shows each participant’s alignment scores. Mean results are shown in the bottom row plots. $\uparrow$ is better.
Figure 6: Participant-model Gromov-Wasserstein distance results for [-pos] (left) and [+pos] (right) conditions. Each subplot shows each participant’s results. $\downarrow$ is better.
@@ 4.5 Brain Scores are Specific to Language
Keeping architecture and learning objective largely constant, substituting natural language with protein folding sequences (i.e., training data) virtually canceled previously observed model-brain correspondences. The non‑linguistic controls obtained brain scores $\leq 0.03$, whereas the poorest text‑trained baseline reached 0.23 ($\approx$8‑fold higher; $\Delta\approx$0.20; Wilcoxon $p<.01$ for every contrast). These results rule out spurious correlations arising from the task itself and demonstrate that exposure to linguistic input is a critical factor for capturing cortical representations of language processing.
@@ 5 Conclusion
The analyzed dataset includes 243 BOLD signals per participant and activations from 19 models. Increasing both the number of subjects and model diversity would increase statistical power and generalizability. Future work should use modality-matched datasets (e.g., spoken-language tasks with audio models, or visual captioning tasks with vision-language models) to disentangle task effects from representational differences.

### v2

@@ 3 Methodology
We used the dataset provided by pereira2018toward to compare brain and language model representations. This is consistent with previous studies (schrimpfNeuralArchitectureLanguage2021a, ; gauthierLinkingArtificialHuman2019, ; schwartzInducingBrainrelevantBias2019, ; ootaNeuralLanguageTaskonomy2022, ; awInstructiontuningAlignsLLMs2024, ) [^1] Dataset available at https://osf.io/crwz7/.
Code: https://github.com/IParraMartin/NCLM-Code
. For the brain model comparisons, we used the fMRI data from 6 adult subjects. The stimuli sentences consisted of simple and natural statements ($N_{\text{stimuli}}=243$), with an average length of 13 words ($\sigma=2.9$). The subjects were indicated to read the sentences naturally and thinking about their meaning. The brain activity was recorded while subjects were presented the sentences one at a time. For each subject, the dataset provides 243 BOLD signals represented as $\approx$200,000-dimensional vectors.
@@ 3.4 Topological Analysis of Brain and LM Layers
In addition to GW, we also computed the centered kernel alignment (CKA) of each participant to each layer, as well as the overall mean for the [+pos] and [-pos] conditions. CKA allowed to measure how similarly two sets of representations encoded relationships across points with minimal intermediate steps. CKA has been previously used in the context of cognitive science cohen1994localization , deep learning kornblith2019similarity and alignment of artificial and biological neural networks murphy2024correcting .
We followed the unbiased kernel alignment formulation kornblith2019similarity (Equation 3). The selection of this variant was motivated by previous findings on the limitations of the original CKA implementation (e.g., murphy2024correcting ). The selected variant has proven robust to high-sample low-dimensionality and low-sample high-dimensionality data setups (murphy2024correcting, ).
$\text{CKA}(K,L)=\frac{\text{HSIC}(K,L)}{\sqrt{\text{HSIC}(K,L)\,\,\text{HSIC}(L,L)}},$ (3)
@@ 4.1 Positional Information is Key for Brain Alignment
Figure 1 shows the mean brain-model correlations for each model class under four conditions: PCA-compressed versus original feature spaces, crossed with positional ([+pos]) and non-positional ([–pos]) internal representations.
@@
Figure 1: A-B: Differences in brain correlations across positional embedding and dimensionality conditions (PCA vs full dimensions) for causal (A) and bidirectional (B) models. C: Selected language processing regions of interest (ROI) of each subject. D-E: Brain-model correlation trends across model depth for bidirectional (D) and causal (E) models.
Layer trajectories in bidirectional models. Patterns diverged across architectures: some displayed a smooth monotonic rising (e.g., distilbert-uncased), whereas others dipped in the middle layers before recovering at the deepest (albert-large-v2) (see Figure 2). Several models maintained above-average [–pos] scores, pointing to implicit positional inference. All bidirectional encoders peaked in the final layers.
@@
Figure 2: All models’ layer-wise brain scores (left bidirectional models, right causal models). The results of the [+pos] condition are shown in blue; orange shows the results for [-pos] condition. Dashed lines show results using PCA data. Purple line indicates overall mean. 0% indicates input layer; 100% corresponds to output layer. Shading indicates range.
@@ 4.2 Scaled Dot-product Attention Accounts for Some Brain Alignment
To understand how models’ internal mechanisms behaved in the observed brain alignment and experimental conditions, we analyzed the scaled dot-product attention heads’ brain scores. Analyzing individualized contributions helped clarify whether the observed brain alignment showed distributed properties or specialization clusters within the attention mechanism.
Multi‑head scaled‑dot product attention is implemented as parallel weight‑matrix operations. Let $W^{Q}_{h}$, $W^{K}_{h}$, $W^{V}_{h}$ $\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ be the query, key, and value projection matrices for head $h\in\{1,\dots,H\}$.
Given a sequence $X\in\mathbb{R}^{n\times d_{\text{model}}}$ we compute
$Q_{h}=XW^{Q}_{h};\quad K_{h}=XW^{K}_{h};\quad V_{h}=XW^{V}_{h}.$
If $\sigma(\cdot)$ denotes a row-wise softmax operation, the attention weights are $\sigma(Q_{h}K_{h}^{\top}/\sqrt{d_{k}})$. Thus, the result for head $h$ is
$\text{head}_{h}=\sigma(\frac{Q_{h}K_{h}^{\top}}{\sqrt{d_{\text{model}}}})V_{h}\in\mathbb{R}^{n\times d_{k}}.$
All $H$ heads are concatenated and projected with $W^{0}\in\mathbb{R}^{Hd_{k}\times d_{\text{model}}}$, producing the final multi-head attention output vaswani2017attention .
Head-wise brain scores were computed as in the layer-wise analysis (§3.3 and §4.1). Figure 3 reports scores for the [-pos] and [+pos] conditions.
@@
Figure 3: Comparison of model-brain alignment across representational levels and architectural configurations. A-B: Average BrainScores for each model under different representational components. The figures show how bidirectional (top) and causal (bottom) architectures differ in their alignment with neural responses. C-D: Empirical cumulative distribution functions (ECDFs) of head-level BrainScores, comparing the distribution of brain alignment across individual attention heads for each model (top bidirectional; bottom causal). Curves positioned further to the right indicate that a greater proportion of heads exhibit higher correlations with brain activity. Across panels, causal models consistently outperform bidirectional models, and the inclusion of positional encoding tends to improve model-brain correspondence, suggesting that positional information enhances the representational similarity between transformer models and neural data.
@@ 4.3 Attention Weights Are Marginally Relevant for Brain Alignment
To account for the impact of raw attention weights on the brain scores we isolated the self-attention operation from the scaled dot-product attention mechanism. To obtain attention weights, the model performs matrix multiplication between the query matrix $Q\in\mathbb{R}^{n\times d_{k}}$ and the transpose of the key matrix $K^{\top}\in\mathbb{R}^{d_{k}\times n}$, which then is scaled by $\sqrt{d_{k}}$ for numerical stability. The resulting similarity matrix is then passed through a softmax so that its entries lie in $[0,1]$. As in previous steps, we computed the brain scores following the same pipeline. The brain scores are shown in Figure 3.
Bidirectional encoders. Across all bidirectional models, attention weights’ brain scores remained uniformly low in the [–pos] and recovered some relevance at the [+pos] condition. However, adding positional encodings produced only marginal gains, far smaller than those observed at layer‑level or scaled dot-product attention analyses. Moreover, the changes did not exhibit a coherent depth‑wise trend: correlations fluctuated idiosyncratically from head to head, suggesting that no specific subset of bidirectional heads was consistently responsible for brain alignment.
@@ 4.4 Brain and Layer Data Spaces Share Characteristics
To analyze the brain-model representational similarities, we computed the layer-wise unbiased CKA [^5] Unbiased CKA is predictable from brain scores. Still, it provides valuable insights. See Appendix B for post-hoc result analysis. and Gromov-Wasserstein distance. The results are provided for each participant and each layer using the most brain aligned causal language model, opt-2.7b.
Layer-wise CKA Analysis. All participants’ brain representations were severely misaligned when the model was deprived of positional information. Interestingly, most subjects showed a similar pattern: CKA steadily decreased to the 0.25-0.37 range (layer 20 to 25) showing a momentous recovery at the 28th layer, and a final decrease from there to the last layer. Adding positional information drastically changed the trends, showing a largely monotonic improvement until reaching the peak similarities ($\approx$0.61-0.87) at the final layers.
@@
Figure 4: Participant-model CKA (top) and Gromov-Wasserstein distance (bottom) results for [-pos] and [+pos] conditions. For GW distance $\downarrow$ is better; for CKA $\uparrow$ is better.
@@ 4.5 Brain Scores are Specific to Language
Figure 5 shows the mean brain scores comparisons for three models trained on non-linguistic tasks and the worst-performing linguistic model (i.e., language model trained on language). All models shared the same masked‑token prediction objective; only the training data differed (protein sequences vs text).
@@
Keeping architecture and learning objective largely constant, substituting natural language with protein folding sequences virtually canceled previously observed model-brain correspondences. The non‑linguistic controls obtained brain scores $\leq 0.03$, whereas the poorest text‑trained baseline reached 0.23 ($\approx$8‑fold higher; $\Delta\approx$0.20; Wilcoxon $p<.01$ for every contrast). These results rule out spurious correlations arising from the task itself and demonstrate that exposure to linguistic input is a critical factor for capturing cortical representations of language processing.
Figure 5: Mean brain scores for non-linguistic control models versus the weakest bidirectional language model, xlm-roberta.
@@ 5 Conclusion
The findings should be interpreted in light of some constraints. First, the analyzed dataset includes 243 BOLD signals per participant and activations from 19 models. Increasing both the number of subjects and model diversity would increase statistical power and generalizability. Second, the head‑wise analysis pooled voxels across cortical regions; no ROI distinctions were drawn. A finer-grained ROI-by-head analysis could reveal differential alignment patterns. Future work should use modality-matched datasets (e.g., spoken-language tasks with audio models, or visual captioning tasks with vision-language models) to disentangle task effects from representational differences.
@@ Appendix B Independence of Empirical Results
Appendix C Data and Code Availability
The code and data used during the experiments are made available here. The GitHub repository includes instruction on how to proceed in order to reproduce the disclosed results.
Code and datasets available upon acceptance.